# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ 8 Robust Lower Bounds for Greedy Trees
To state this more formally, we first note that the Fourier graph of $f^{*}_{-\mathcal{B}}$ is equivalent to the induced subgraph of $\𝒮\mathcal{S}\backslash\mathcal{B}$ in $\mathcal{G}_{f^{*}}$, denoted $\𝒮\mathcal{G}_{f^{*}}[\mathcal{S}\backslash\mathcal{B}]$. Denote the connected component of the empty set $\emptyset$ in $\𝒮\mathcal{G}_{f^{*}}[\mathcal{S}\backslash\mathcal{B}]$ as $\mathcal{S}_{-\mathcal{B},\textsf{MSP}}$ and set $S^{*}_{-\mathcal{B},\textsf{MSP}}\coloneqq\cup_{S\in\mathcal{S}_{-\mathcal{B},\textsf{MSP}}}S$. Furthermore, let $\mathcal{S}_{-\mathcal{B},-\textsf{MSP}}\coloneqq\mathcal{S}\backslash\left(\mathcal{S}_{-\mathcal{B},\textsf{MSP}}\cup\mathcal{B}\right)$ denote the subcollection comprising all vertices disconnected from $\emptyset$ in $\𝒮\mathcal{G}_{f^{*}}[\mathcal{S}\backslash\mathcal{B}]$. We have the following elaboration on Theorem 5.1.
@@
Under Assumption 3.1, suppose $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}\sim_{\text{i.i.d.}}\mathcal{N}(0,\sigma^{2})$, and let $\mathcal{G}_{f_{0}^{*}}$ be the Fourier graph of $f_{0}^{*}$ with weight function $w$. Let $\mathcal{B}\subset\mathcal{S}$ be a subset of vertices of $\mathcal{G}_{f_{0}^{*}}$. Let $s_{\operatorname{T}}$ be the minimum size of a traversal $T$ for $\mathcal{S}_{-\mathcal{B},-\textsf{MSP}}$ such that $T\subset[d]\backslash S^{*}_{-\mathcal{B},\textsf{MSP}}$ and let $s_{-\mathcal{B},\textsf{MSP}}\coloneqq\left|S^{*}_{-\mathcal{B},\textsf{MSP}}\right|$. For any $10<\delta<$ 1, suppose
@@ 9 Extensions of Lower Bounds
Similar to discrete covariate setting, there is marginal signal bottleneck, and we may completely imitate the proof of Lemma 4.2, except that we can no longer use the results from Appendix B to bound the query path length. We can enforce this by constraining the algorithm to make relatively balanced splits on irrelevant covariates (i.e., whenever $\mathcal{C}_{L}$ and $\mathcal{C}_{R}$ are the children resulting from a chosen split, the log ratio $\log\left(\nu(\mathcal{C}_{L})/\nu(\mathcal{C}_{R})\right)$ has to be bounded with high probability). With this, we can obtain an estimation error lower bound for the CART and RF estimators whenever $\log_{2}n\leq Cd$ for some universal $C$.
@@
To see how far we can push this result, we next construct the equivalent of the Fourier graph. To this end, we first index each basis element using a collection of tuples, i.e., given $\Lambda\coloneqq\left\{(k_{j},l_{j},m_{j})\colon j=1,2,\ldots,r\right\}$, define $\chi_{\Lambda}(\mathbf{x})\coloneqq\prod_{j=1}^{r}\psi_{l_{j},m_{j}}(x_{k_{j}})$. Now given any function $f^{*}\in L^{2}\left([0,1]^{d}\right)$, we may write it as $f^{*}\coloneqq\sum_{\Lambda\in\mathfrak{L}}\alpha_{\Lambda}\chi_{\lambda}$. Define the directed hypergraph $\mathcal{G}$ with vertex set $\mathfrak{L}$ and put an edge $\left\{\Lambda^{(1)},\Lambda^{(2)},\ldots,\Lambda^{(q)}\right\}\to\Lambda^{(q+1)}$ if the following hold: (i) the union $\cup_{j=1}^{q}\Lambda^{(j)}$ contains at most one entry for each covariate, (ii) $\left|\Lambda^{(q+1)}\backslash\cup_{j=1}^{q}\Lambda^{(j)}\right|\leq 1$, and (iii) if $\Lambda^{(q+1)}\backslash\cup_{j=1}^{q}\Lambda^{(j)}=\left\{(k,l,m)\right\}$, then $(k,l-1,\lceil m/2\rceil)\in\cup_{j=1}^{q}\Lambda^{(j)}$. Similar to the binary case, we define a weight function via $w(\Lambda)\coloneqq\alpha_{\Lambda}^{2}$ and call $(\mathcal{G},w)$ the wavelet graph of $f^{*}$.
@@ 10 Overcoming Marginal Signal Bottleneck
Within the family of greedy algorithms, Loh (2002)’s GUIDE algorithm tests for pairwise interactions when selecting splits and is thus able to estimate the XOR function. On the other hand, it is unable to estimate higher-order monomials. CART can be modified to allow for splits that are oblique, that is, along linear hyperplanes depending on multiple covariates (Cattaneo et al., 2024). This creates fewer opportunities for signal bottleneck—for instance, the XOR function in continuous covariates is no longer an issue. However, optimizing over oblique splits is vastly more computationally intensive and also reduces the interpretability of the resulting model.
@@ Acknowledgments
Appendix A Lower Bounds for Non-Adaptive Trees
We say that a regression tree algorithm $\mathcal{A}$ is non-adaptive if $\mathcal{A}_{t}$ does not depend on the response data $Y_{1},Y_{2},\ldots,Y_{n}$ (see Section 2). We call the output of such an algorithm a non-adaptive tree model, denoting it using $\hat{f}_{\operatorname{NA}}(-;\mathcal{D}_{n},\Theta)$. This class of tree models include completely random trees, which are grown iteratively, similar to CART, except that at each iteration, the split feature for a new cell $\mathcal{C}$ is chosen uniformly at random from $[d]\backslash J(\mathcal{C})$. This can also be generalized to incorporate non-uniform split probabilities. Another subclass comprises Mondrian trees (Roy and Teh, 2008), which are defined via a recursive generative process that selects splits according to the length of each side of a cell and uses a stochastic stopping condition. These tree models are more appropriately defined over continuous covariates. As is the case with greedy trees, non-adaptive trees can be combined in ensembles, with the resulting estimators described as some type of random forests (Lakshminarayanan et al., 2014; Biau, 2012). We will denote an ensemble of non-adaptive tree models using $\hat{f}_{\operatorname{NA-E}}(-;\mathcal{D}_{n},\Theta)$.
@@ Appendix A Lower Bounds for Non-Adaptive Trees
$\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)\coloneqq\mathbb{E}_{S^{*},\mathcal{D}_{n},\Theta}\left\{R\left(\hat{f}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\right\}.$ (A.1)
@@
Lemma A.2.
Let $\hat{f}$ be a regression tree ensemble model in which each tree is fit using a regression tree algorithm $\mathcal{A}$. Then under Assumption A.1, its expected $L^{2}$ risk satisfies $\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)\geq\left(1-\kappa\delta^{1/2}\right)\operatorname{Var}\left\{f^{*}_{0}(\mathbf{X})\right\}$, where $\delta\coloneqq\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}$ and $\kappa\coloneqq 2\|f_{0}^{*}\|_{\infty}/\operatorname{Var}\left\{f^{*}_{0}(\mathbf{X})\right\}^{1/2}$.
@@
$\begin{split}\mathbb{E}_{\mathbf{X},S^{*}}\left\{\operatorname{Var}_{\mathbf{X}}\left\{h(\mathbf{X})~{}|~{}\mathbf{X}_{[d]\backslash S^{*}}=\mathbf{x}_{[d]\backslash S^{*}}\right\}^{1/2}\right\}&\leq\mathbb{E}_{\mathbf{X},S^{*}}\left\{h(\mathbf{X})^{2}\right\}^{1/2}\\ &\leq\|f_{0}^{*}\|_{\infty}\mathbb{P}_{\mathbf{X},\mathcal{D}_{n},\Theta,S^{*}}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}^{1/2}\\ &=\|f_{0}^{*}\|_{\infty}\delta^{1/2}.\end{split}$ (A.2)
@@
Lemma A.3.
Suppose $\mathcal{A}_{t}$ is non-adaptive. Then under Assumption A.1, the probability of covariate selection is upper bounded as $\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}\leq(s/d)\log_{2}n$.
@@
$\begin{split}\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}&=\mathbb{E}\left\{\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset~{}|~{}\mathbf{X},\mathcal{D}_{n},\Theta\right\}\right\}.\end{split}$ (A.3)
@@
$\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset~{}|~{}\mathbf{X},\mathcal{D}_{n},\Theta\right\}\leq\frac{s\left|J(\mathbf{X};\mathcal{D}_{n},\Theta)\right|}{d}.$ (A.4)
Let $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{K}$ be the leaves in $\mathfrak{T}$ (still conditioning on $\mathcal{D}_{n}$ and $\Theta$). Notice that $\left|J(\mathbf{X};\mathcal{D}_{n},\Theta)\right|$ is the length of the query path of $\mathbf{X}$ and is equal to $-\log_{2}\nu(\mathcal{L}_{k})$ if $\mathbf{x}\in\mathcal{L}_{k}$. We therefore have
$\mathbb{E}\left\{\left|J(\mathbf{X};\mathcal{D}_{n},\Theta)\right|~{}|~{}\mathcal{D}_{n},\Theta\right\}=-\sum_{k=1}^{K}\nu(\mathcal{L}_{k})\log_{2}(\nu(\mathcal{L}_{k})),$ (A.5)
@@
Proposition A.4.
Under Assumption A.1, for any $10<\delta<$ 1, suppose $n\leq 2^{\delta d/s}$. Then the expected $L^{2}$ risk of any non-adaptive tree model satisfies
$\mathfrak{R}\left(\hat{f}_{\operatorname{NA}},f^{*}_{0},d,n\right)\geq(1-\delta)\operatorname{Var}\left\{f^{*}(\mathbf{x})\right\}.$ (A.6)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{NA-E}},f^{*}_{0},d,n\right)\geq\left(1-\kappa\delta^{1/2}\right)\operatorname{Var}\left\{f^{*}(\mathbf{x})\right\},$ (A.7)
@@
Equation (A.6) follows immediately from Lemma 3.2 and Lemma A.3. Equation (A.7) follows similarly using Lemma A.2 and Lemma A.3. ∎
Appendix B Length of Query Paths
Let $\mathcal{Q}(\mathbf{x};\mathcal{D}_{n},\Theta)$ denote the query path of a tree fitted using a greedy tree algorithm to a dataset $\mathcal{D}_{n}=\left\{(\mathbf{X}_{i},Y_{i})\colon i=1,2,\ldots,n\right\}$ with a uniform distribution on the covariate space {±$\{\pm 1\}^{d}$. Assume also that the responses $Y_{1},Y_{2},\ldots,Y_{n}$ are completely independent of the covariates. We let $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{l}$ denote the cells in $\mathcal{Q}(\mathbf{x};\mathcal{D}_{n},\Theta)$, and further denote $n_{t}\coloneqq N(\mathcal{C}_{t})$ for $t=0,1,\ldots,l$. Observe also that $l=\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|$.
Lemma B.1.

### v3

@@ 8 Robust Lower Bounds for Greedy Trees
To state this more formally, we first note that the Fourier graph of $f^{*}_{-\mathcal{B}}$ is equivalent to the induced subgraph of $\mathcal{S}\backslash\mathcal{B}$ in $\mathcal{G}_{f^{*}}$, denoted $\mathcal{G}_{f^{*}}[\mathcal{S}\backslash\mathcal{B}$]. Denote the connected component of the empty set $\emptyset$ in $\mathcal{G}_{f^{*}}[\mathcal{S}\backslash\mathcal{B}$] as $\mathcal{S}_{-\mathcal{B},\textsf{MSP}}$ and set $S^{*}_{-\mathcal{B},\textsf{MSP}}\coloneqq\cup_{S\in\mathcal{S}_{-\mathcal{B},\textsf{MSP}}}S$. Furthermore, let $\mathcal{S}_{-\mathcal{B},-\textsf{MSP}}\coloneqq\mathcal{S}\backslash\left(\mathcal{S}_{-\mathcal{B},\textsf{MSP}}\cup\mathcal{B}\right)$ denote the subcollection comprising all vertices disconnected from $\emptyset$ in $\mathcal{G}_{f^{*}}[\mathcal{S}\backslash\mathcal{B}$]. We have the following elaboration on Theorem 5.1.
@@
Under Assumption 3.1, suppose $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}\sim_{\text{i.i.d.}}\mathcal{N}(0,\sigma^{2})$, and let $\mathcal{G}_{f_{0}^{*}}$ be the Fourier graph of $f_{0}^{*}$ with weight function $w$. Let $\mathcal{B}\subset\mathcal{S}$ be a subset of vertices of $\mathcal{G}_{f_{0}^{*}}$. Let $s_{\operatorname{T}}$ be the minimum size of a traversal $T$ for $\mathcal{S}_{-\mathcal{B},-\textsf{MSP}}$ such that $T\subset[d]\backslash S^{*}_{-\mathcal{B},\textsf{MSP}}$ and let $s_{-\mathcal{B},\textsf{MSP}}\coloneqq\left|S^{*}_{-\mathcal{B},\textsf{MSP}}\right|$. For any $0<\delta<1$, suppose
@@ 9 Extensions of Lower Bounds
Similar to discrete covariate setting, there is marginal signal bottleneck, and we may completely imitate the proof of Lemma 4.2, except that we can no longer use the results from Appendix C to bound the query path length. We can enforce this by constraining the algorithm to make relatively balanced splits on irrelevant covariates (i.e., whenever $\mathcal{C}_{L}$ and $\mathcal{C}_{R}$ are the children resulting from a chosen split, the log ratio $\log\left(\nu(\mathcal{C}_{L})/\nu(\mathcal{C}_{R})\right)$ has to be bounded with high probability). With this, we can obtain an estimation error lower bound for the CART and RF estimators whenever $\log_{2}n\leq Cd$ for some universal $C$.
@@
[…]
@@
Finally, we note that while our current analysis is tailored to the uniform distribution and relies on the structure of Fourier or wavelet bases, recent work in computational learning theory has pursued more distribution-agnostic complexity characterizations. For example, Joshi et al. (2024) analyze the statistical and gradient query complexity of learning sparse functions under general product measures.
@@ 10 Overcoming Marginal Signal Bottleneck
Within the family of greedy algorithms, Loh (2002)’s GUIDE algorithm tests for pairwise interactions when selecting splits and is thus able to estimate the XOR function. On the other hand, it is unable to estimate higher-order monomials. One can generalize this idea in a binary feature setting to permit splits on $k$-way monomials $\prod_{j\in S}x_{j}$ with $|S|\leq k$. Such an algorithm can indeed learn leap-$k$ regression functions (e.g. order $k$ monomials, see also Appendix K), under a genericity condition on Fourier coefficients, with a much lower sample complexity of $O(k\log d)$, but at a much higher computational cost of $O(nd^{k})$ per split. However, for functions beyond leap-$k$ (e.g., monomials of order larger than $k$), all conditional correlations of degree $\leq k$ vanish, so consistency still requires $\exp(\Omega(d))$ samples. Alternatively, CART can be modified to allow for splits that are oblique, that is, along linear hyperplanes $\sum_{j=1}^{d}\beta_{j}x_{j}$ depending on multiple covariates (Cattaneo et al., 2024). This creates fewer opportunities for signal bottleneck—for instance, the XOR function in continuous covariates is no longer an issue. However, optimizing over oblique splits is vastly more computationally intensive and also reduces the interpretability of the resulting model.
@@
11 Simulations
To supplement our theory, we perform a simulation study in which we generate training data according to (1), with the covariates distributed uniformly on $\{\pm 1\}^{d}$ and the regression function chosen to be $f^{*}(\mathbf{x})=x_{1}x_{2}+\alpha x_{1}$. This corresponds to the XOR function when $\alpha=0$, while setting nonzero $\alpha$ allows for some amount of marginal signal. We vary $d\in\{10,20,30,40,50\}$, $\log_{2}n\in\{7,9,11,13,15\}$, $\alpha\in\{0,0.02\}$, and $\sigma^{2}\in\{0,0.1\}$. We use two metrics to evaluate the performance of CART on each dataset:
  • Expected mean squared error (MSE) $R(\hat{f}_{\operatorname{CART}},f^{*})$, where $\hat{f}_{\operatorname{CART}}$ is fitted with a minimum impurity decrease stopping criterion (14). Here, the threshold $\gamma$ is selected based on a held-out validation set.
  • Split coverage of each feature $X_{k}$ by a CART tree $\mathfrak{T}$ that is grown to purity, that is, with no stopping criterion. Split coverage is defined as:
    $\text{coverage}(k,\mathfrak{T})\coloneqq\sum_{\mathcal{L}\in\mathcal{P}(\mathfrak{T})}2^{-\text{depth}(\mathcal{L})}\mathbb{1}\left\{\text{path from root to}~\mathcal{L}~\text{splits on}~X_{k}\right\}.$ (57)
    For the uniform measure, $\text{coverage}(k,\mathfrak{T})$ computes the probability that the root-to-leaf path in $\mathfrak{T}$ followed by a random query point $\mathbf{X}$ splits on the feature $X_{k}$.
Selected results from the experiment, averaged over 200 replicates, are shown in Figure 1 and Figure 2. In each panel of Figure 1, we plot how expected MSE varies across the grid of $n$ and $d$ values, stratifying on all other hyperparameters of the data generating process. In both panels, we focus on the noiseless setting ($\sigma^{2}=0$), under which the null model satisfies $R(0,f^{*})=1+\alpha^{2}$. The left panel corresponds to the XOR setting ($\alpha=0$), while the right panel corresponds to $\alpha=0.02$. Note that while $d$ varies on a linear scale, $n$ varies on a logarithmic scale. Nonetheless, in the left panel, we see that the expected MSE values increase along the bottom-left to top-right diagonal, reflecting that the sample complexity required to “learn” the regression function has exponential dependence on the ambient dimension, which is in agreement with Proposition 4.1. This exponential dependence disappears in the right panel, even though the energy of the additional term, $0.02X_{1}$, is much smaller than that of the main term $X_{1}X_{2}$. This agrees with the high-dimensional consistency guarantee provided in Theorem 7.5. Figure 2 similarly plots how split coverage varies across the grid of $n$ and $d$ values, fixing $\alpha=0$ and $\sigma^{2}=0$. For this figure, however, the left panel corresponds to the coverage of $X_{2}$, a relevant feature, while the right panel corresponds to the coverage of $X_{3}$, which is an irrelevant feature. In the left panel, we see that coverage of $X_{2}$ has a decreasing trend across the bottom-left to top-right diagonal, which again shows that the sample complexity required for reliably performing feature selection grows exponentially in $d$, in agreement with Lemma 4.2. The right panel gives reference values for coverage of an irrelevant feature. One may observe that these values are upper bounded by $\log_{2}n/d$. Additional simulation results, including those for $\sigma^{2}=0.01$, are shown in Appendix L.
Figure 1: Expected MSE $R(\hat{f}_{\operatorname{CART}},f)$ for CART fitted to $Y=X_{1}X_{2}+\alpha X_{1}$, with $\alpha=0$ (left panel) and $\alpha=0.02$ (right panel). CART is fitted with a minimum impurity decrease stopping criterion. We observe high-dimensional consistency in the right panel since the true function satisfies MSPand lack of high-dimensional consistency in the left panel since the true function does not satisfy MSP.
Figure 2: Split coverage for features $X_{2}$ and $X_{3}$ for CART fitted to $Y=X_{1}X_{2}$, without any stopping criterion. Split coverage is defined in (57). We observe that the sample size required to reliably split on the relevant feature grows exponentially in $d$.
@@ Acknowledgments
Appendix A Consistency of ERM trees, Greedy Trees, and Non-Adaptive Trees
Machine learning models are often fit by solving empirical risk minimization (ERM). Solutions to ERM over appropriate classes of decision tree functions can be shown to be consistent, but unfortunately are computationally infeasible to obtain. For instance, ERM for classification trees is known to be NP-complete in the worst case (Hyafil and Rivest, 1976) when the attribute count $d$ is part of the input. Their reduction depends only on the choice of split variables and tree topology, so the same hardness applies verbatim to the squared‐error regression objective studied in this work. If $d$ is fixed, the input space has only $2^{d}$ patterns, so one can enumerate the $\leq d^{2^{s}}$ depth-$s$ trees and evaluate each in $O(n)$ time, yielding a (large-constant) polynomial algorithm; hardness disappears in this constant-$d$ corner case (Donoho, 1997; Chatterjee and Goswami, 2021). Whereas, our focus in this work is on the high-dimensional regime ($d\to\infty$, $\log n/d\to 0$) where exhaustive ERM needs $O(nd^{2^{s}})$ operations while CART fits a depth-$s$ tree in $O(nd)$.
Practitioners make use of greedy algorithms such as CART  (Breiman et al., 1984), ID3 (Quinlan, 1986), or C4.5 (Quinlan, 1993). Such algorithms grow a tree in a top-down manner, with the split for a given node chosen by optimizing a local node-specific objective. Once made, a split is never reversed. [^6] Decision trees are sometimes pruned as a one-off post-processing procedure after they are fully grown. These heuristics seem to work well in practice—both RFs and gradient boosting are built out of CART trees—but are notoriously difficult to analyze. Indeed, early efforts to prove consistency of greedy regression and classification trees under standard smoothness conditions on $f^{*}$ had to further assume that the bias of the tree converged to zero (Breiman et al., 1984; Chaudhuri et al., 1995). More recent works were able to replace this assumption with more reasonable alternatives on $f^{*}$ such as additivity  (Scornet et al., 2015; Klusowski, 2020; Klusowski and Tian, 2024) or a submodularity condition (Syrgkanis and Zampetakis, 2020). Chi et al. (2022) further introduced and derived consistency under the sufficient impurity decrease (SID) condition (7). Mazumder and Wang (2024) later improved upon the consistency rates of Chi et al. (2022) and showed that the SID condition is satisfied for larger classes of models.
In another line of work, researchers have studied decision trees whose splits are non-adaptive, meaning that the partitioning is chosen independently of the response $Y$. Examples include purely random-split trees (Biau et al., 2008; Biau, 2012; Klusowski, 2021), Mondrian trees which split according to a Bayesian model of the covariate distribution (Roy and Teh, 2008; Lakshminarayanan et al., 2014; Mourtada et al., 2017; Cattaneo et al., 2023), and $k$-Potential Nearest Neighbors (Lin and Jeon, 2006; Biau and Devroye, 2010; Scornet, 2016; Biau and Scornet, 2016; Shi et al., 2024). These methods are valuable as theoretical baselines but are generally not used in practice. In particular, purely random forests fall in this class, whereas methods such as honest trees remain adaptive in their split selection and thus do not. Despite having consistency guarantees under fewer assumptions, non-adaptive trees typically underperform greedy, label-driven procedures in terms of adaptivity to heterogeneous truth.
Appendix B Lower Bounds for Non-Adaptive Trees
[…]
@@ Appendix B Lower Bounds for Non-Adaptive Trees
$\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)\coloneqq\mathbb{E}_{S^{*},\mathcal{D}_{n},\Theta}\left\{R\left(\hat{f}(-;\mathcal{D}_{n},\Theta),f^{*}\right)\right\}.$ (B.1)
@@
Lemma B.2.
Let $\hat{f}$ be a regression tree ensemble model in which each tree is fit using a regression tree algorithm $\mathcal{A}$. Then under Assumption B.1, its expected $L^{2}$ risk satisfies $\mathfrak{R}\left(\hat{f},f^{*}_{0},d,n\right)\geq\left(1-\kappa\delta^{1/2}\right)\operatorname{Var}\left\{f^{*}_{0}(\mathbf{X})\right\}$, where $\delta\coloneqq\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}$ and $\kappa\coloneqq 2\|f_{0}^{*}\|_{\infty}/\operatorname{Var}\left\{f^{*}_{0}(\mathbf{X})\right\}^{1/2}$.
@@
[…]
@@
Lemma B.3.
Suppose $\mathcal{A}_{t}$ is non-adaptive. Then under Assumption B.1, the probability of covariate selection is upper bounded as $\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}\leq(s/d)\log_{2}n$.
@@
$\begin{split}\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset\right\}&=\mathbb{E}\left\{\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset~|~\mathbf{X},\mathcal{D}_{n},\Theta\right\}\right\}.\end{split}$ (B.3)
@@
$\mathbb{P}\left\{J(\mathbf{X};\mathcal{D}_{n},\Theta)\cap S^{*}\neq\emptyset~|~\mathbf{X},\mathcal{D}_{n},\Theta\right\}\leq\frac{s\left|J(\mathbf{X};\mathcal{D}_{n},\Theta)\right|}{d}.$ (B.4)
Let $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{K}$ be the leaves in $\mathfrak{T}$ (still conditioning on $\mathcal{D}_{n}$ and $\Theta$). Notice that $\left|J(\mathbf{X};\mathcal{D}_{n},\Theta)\right|$ is the length of the query path of $\mathbf{X}$ and is equal to $\log_{2}\nu(\mathcal{L}_{k})$ if $\mathbf{x}\in\mathcal{L}_{k}$. We therefore have
$\mathbb{E}\left\{\left|J(\mathbf{X};\mathcal{D}_{n},\Theta)\right|~|~\mathcal{D}_{n},\Theta\right\}=-\sum_{k=1}^{K}\nu(\mathcal{L}_{k})\log_{2}(\nu(\mathcal{L}_{k})),$ (B.5)
@@
Proposition B.4.
Under Assumption B.1, for any $0<\delta<1$, suppose $n\leq 2^{\delta d/s}$. Then the expected $L^{2}$ risk of any non-adaptive tree model satisfies
$\mathfrak{R}\left(\hat{f}_{\operatorname{NA}},f^{*}_{0},d,n\right)\geq(1-\delta)\operatorname{Var}\left\{f^{*}(\mathbf{x})\right\}.$ (B.6)
@@
$\mathfrak{R}\left(\hat{f}_{\operatorname{NA-E}},f^{*}_{0},d,n\right)\geq\left(1-\kappa\delta^{1/2}\right)\operatorname{Var}\left\{f^{*}(\mathbf{x})\right\},$ (B.7)
@@
Equation (B.6) follows immediately from Lemma 3.2 and Lemma B.3. Equation (B.7) follows similarly using Lemma B.2 and Lemma B.3. ∎
Appendix C Length of Query Paths
Let $\mathcal{Q}(\mathbf{x};\mathcal{D}_{n},\Theta)$ denote the query path of a tree fitted using a greedy tree algorithm to a dataset $\mathcal{D}_{n}=\left\{(\mathbf{X}_{i},Y_{i})\colon i=1,2,\ldots,n\right\}$ with a uniform distribution on the covariate space $\{\pm 1\}^{d}$. Assume also that the responses $Y_{1},Y_{2},\ldots,Y_{n}$ are completely independent of the covariates. We let $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{l}$ denote the cells in $\mathcal{Q}(\mathbf{x};\mathcal{D}_{n},\Theta)$, and further denote $n_{t}\coloneqq N(\mathcal{C}_{t})$ for $t=0,1,\ldots,l$. Observe also that $l=\left|J(\mathbf{x};\mathcal{D}_{n},\Theta)\right|$.
Lemma C.1.